### v1

@@ 2.3 One segment
For the discussion below, we assume that the segment has a two-dimensional surface boundary, and we discuss the gradients along a line, perpendicular to that plane’s surface. We neglect the electrokinetic and surface-related electric effects connected to membranes and nanopores [17] and focus on the balanced state of membranes in ionic solutions. We compose the segments from such layers (sheets), $(x,y)$ parallel plates, and describe the gradients in the direction of $z$. We introduce the concept of a ’thin physical layer’ parallel to the membrane, characterized by a finite width.
@@
where $N_{A}$ is the number of ions in a $millimol$ and $e_{el}$ is the elementary charge. Notice that all ions contribute to the charge density, so simply $c=\sum_{k}C_{k}$.
@@ 2.4 Two segments
The behavior near the boundary layers is problematic, mainly because the layers themselves are ill-defined. Physically, they are not ideally plain and well-defined surfaces; the layer thicknesses are limited by the size of the ions/molecules and their associated ions; the layers do not necessarily satisfy the ideal requirements of the respective physical laws. Mathematically, a single ’universal’ function that spans all regions surely cannot accurately describe the behavior of the physical model. So, we chose a piecewise description where the pieces seamlessly fit at the boundaries. We assume that between the surfaces of the membrane, classical electricity describes the electric field; the charged layer and the electrolyte jointly produce the electric field; we have a uniformly charged layer from separated charges on the surface and a polarized electrolyte outside them.
Unlike in the classical theory, the electrolyte segment next to the surface layer contains dipole molecules that have more or less balanced charges (the polarization also depends on the external electric field), so they have much less mobility than the dissociated ions: their size and mass are bigger, and their electric force is a fragment of that of the ions. Their virtual charge becomes apparent only in the appropriate environment, yet it creates an electric field similar to a real charge. The final reasons for creating virtual charges are real charges and/or external fields, so an electric field due to virtual charges can also exist inside dielectric matter. We assume the membrane is transparent for the electrical interaction (the electrical field affects the ions in the other segment on the other side of the membrane) but not for their masses (mechanically separates the segments). The membrane does not need to generate any counterforce (except for collisions).
@@ 2.4.1 Charge layers in segments
We consider that the segment is composed of electrically conducting discs (the ions are free to move on the surface) and the charged discs’ contribution to the electric field at point $P$ (see Fig.  2) can be calculated as known from the theory of electricity. Due to the symmetry in the $z$ direction, in a homogeneous solution, the resulting electric field in the plane perpendicular to the $z$ direction is zero, as we have contributions of equal magnitude with opposite signs.
@@ 2.4.2 Simple conducting sheet (a physical condenser)
First, we consider a condenser that obeys the laws of classical electricity and has the geometrical size of a biological neuron. We apply the physics terms to that pseudo-biological object and test whether the values we derive are consistent with physiological phenomena. A tiny (according to [25], page 12, about $2*10^{-3}$) portion of the positive ions leaves their negative counterions behind and forms a thin positively charged ion layer on the surface of the membrane (of course, the same happens on the opposite side with negative ions). Experience shows that – also in biology – two very thin layers of ions are formed on the two sides of the membrane of finite width, see the caption of figure (Fig. 11.22 in [26]): "The ions that give rise to the membrane potential lie in a thin ($<1\ nm$) surface layer close to the membrane" (As shown below, these ions give rise only to part of the potential, and the dipoles in the electrolyte contribute the rest). The two surfaces are covered by a sub-nanometer-thick conducting layer on the top of a $5~{}nm$ thick insulator. So, we model the cell with two finite-width "conductive sheet" layers. These layers represent an electric condenser (so we can calculate the internal electric field between the plates) and counterbalance each other’s electric field outside the condenser. We may assume that the solid surface represents the needed counterforce to keep the charges at rest.
We cannot expect geometrically flat surfaces, as some biological objects of up to $1~{}nm$ in size are present on the surface. So, we must assume a physical ’uniformly charged sheet’ with an average thickness $\Delta z=0.5~{}nm$ on the surface. For the sake of simplicity, we assume a step-like function for the electric field. It is zero in the segment outside the "conducting sheets" (the ’bulk’ portions) and inside the sheets. Between the sheets, it jumps to the value of the electric field of a charged sheet with surface density $\sigma_{A}$ (see Eq.(9)). (In this picture, we see a jump in the value of the electric field on the two sides of the physical layer. We assume that the physical layer somehow emulates the well-defined boundary on the side opposite to that proximal to the membrane. Here, we still assume that a mechanical counterforce due to the surface’s roughness keeps the charges inside the charged layer.)
We assume that all unbalanced (dissociated) ions are in the layer, and the rest contains no dissociated ions; furthermore, the concentration of the ions is unchanged in the segment. In the case of a $400~{}mM$ solution and a $0.5~{}nm$ ’thin physical layer’ Eq.(10) evaluates to an electric field
@@
and at a $5\ nm$ membrane thickness, the voltage across the plates becomes
@@
Our simple model seems to be a strong oversimplification. The number of uncompensated ions needed for the cell, using the method in [25], is $0.88*10^{7}$. The calculated values are about a factor of $5$ lower than the experimentally derived values. "An electrical potential difference about $50-100\ mV$ exists across a plasma membrane only about $5\ nm$ thick, so that the resulting voltage gradient is about $100,000\ V/cm$[26]. "The number of uncompensated ions needed for the cell is $4.7*10^{7}$[25]. We are in the correct order of magnitude, but we arbitrarily assumed a layer thickness, a non-dielectric medium, and a physically unreasonable step-like electric field. The deviation from the experienced values suggests that dipoles in the electrolyte bulk significantly change the achievable electric field and potential across the plates.
Figure 3: The oppositely charged parallel surfaces of the membrane are treated like conducting plates (infinite planes, neglecting fringing), then Gauss’ law can be used to calculate the electric field between the plates. Presuming the plates are at equilibrium with zero electric fields inside the conductors, the result from a charged conducting surface can be used. “from Hyperphysics by Rod Nave, Georgia State University". ©http://hyperphysics.phy-astr.gsu.edu/hbase/
A usual parallel plate condenser (shown in Fig. 3) can be derived with such a model. We have two infinite conducting sheets and an insulator layer between them. The magnitude of the opposite charges on the two plates must be the same as in the classical picture. Given that the amount of surface charge densities is concluded from the ionic concentration in the segment (in a balanced state), the sum of ionic concentrations must be the same in the segments. A constant electric field is present inside the membrane (across the plates of the condenser), and there is zero electric field inside the parallel conducting plates and outside the condenser. In this ideal picture, the charges are aligned on the border of two (infinitely thin) conducting layers and cannot move. The attractive force between them and the opposite charges on the other plate keeps them fixed in the direction of $z$, and the repulsive force between the charges with the same sign keeps their surface density in the $(x,y)$ plane uniform: the infinitely thin plates are equipotential. This picture illustrates the equilibrium state of charges, consisting of infinitely small non-dissociating charge carriers and perfectly smooth surfaces.
@@ 2.4.3 Condenser in dielectric material
When explaining the effect of dielectricity, we must explicitly consider the duality of ions that they obey laws of electricity and thermodynamics simultaneously; furthermore, the complexity of the electric structure of the solution and that the charge carriers have finite size (see "electron size" vs. "dipole size"). According to the theory of electricity, the free charge carriers (the dissociated ions) are located on the surfaces where they form powerfully charged thin layers (condenser’s plates) in the segments separated by a membrane on the two surfaces of the membrane; two proximal charge layers on the surfaces of the membrane. We assume that those ions behave as point charges, i.e., due to the attraction from the opposite charges on the opposite plate, they do not produce an electric field toward the side of the bulk of the electrolyte layers. In the classical picture, those layers represent step-like gradients in the electric field along the $z$ axis, and the constraint that the ions must not enter the membrane provides the needed counterforce.
@@
From a thermodynamical point of view, a driving force acts as long as a concentration gradient between neighboring layers exists. From an electrical point of view, the layers are simple parallel-plate condensers that produce no electric field outside their closed volume and are connected serially; plus, the top layer has free ions. The electric field is proportional to the bulk concentration in the layer. Changes in the local electrical gradient may also alter the degree of dissociation and polarization, producing a graded local electric field. These two driving forces have opposite directions and in a balanced state, the same magnitude. The counterforce acts in a way that constraints the molecules to stay in their layers; it adapts (also compensates for the different interaction speeds) while the gradients are changing. It presses the proximal layer to the membrane and the neighboring layers to each other. (In other words, the internal electric force due to the forcefully changed polarization continuously decreases the electric field and so the potential gradient and so the concentration gradient.)
@@
Due to the membrane’s finite width and surface roughness, the conducting layer has a finite width. The electric field is step-like, with a step size equal to the size of dipoles; furthermore, it is homogeneous in the layers. The rest of that electric field toward the bulk layer directs the dipoles proximal to the neighboring dipole layer, and their polarization increases; in other words, they "produce" an electric field. The final effect is that a dipole layer is attracted to the charged layer, and it forms another layer that shows a charged layer toward the bulk. The process repeats and results in a decaying electric field based on the distance from the membrane’s surface.
@@
In our model, we build the volume of the electrolyte from thin layers of directed dipole molecules (the size of molecules limits the thinness $\Delta z$) in the volume separated into two segments with electrolytes by a membrane with a finite thickness $d$ (that is, a point’s distance in the electrolyte from the two disks will be $z$ and $z+d$, respectively). The contributing potential of an infinitesimal volume on the axis at point $z$ due to the charged sheet is
@@
[…]
@@ 2.5 Bridging the micro- and macro views
We assumed that the function can be interpreted for regions $(-\infty,0)$ and $(1,+\infty)$. However, we have arrived at the boundary between quantized and continuous electricity. The minimal thickness of layers is limited, at least by the size of the ion. Furthermore, the membrane’s surface is not flat; there are structures (mainly lipids) with sizes of up to $1~{}nm$, so it is likely realistic to consider a layer thickness of $0.5~{}nm$ as we did above (on our mathematical scale $0.1=\frac{0.5~{}nm}{5~{}nm}$), for which the multiplier is $1.7$. A similar calculation using a layer thickness of $0.1~{}nm$ (on our mathematical scale, 0.02) results in a multiplier of $3.2$; so we use a multiplier of value 3 (that coresponds to a layer thickness of $0.125~{}nm$ (the radius of a $Na^{+}$ ion is $0.116~{}nm$)). Correspondingly, we assume an electric field
@@
due to the dielectricity in the segment. This contribution is from the dielectricity in the segment. It is to be added to the value of the classic contribution, the field generated by the conducting plates, so the final multiplier is $4$. (The experience is known in technical electricity: the electrolytic capacitors achieve several times higher charge storage capacity by using (pseudocapacitance). By using "roughened anode foil", the thickness of the electrolyte is increased, and the roughening provides the necessary mechanical support. A thicker electrolyte layer wraps the condenser plates, and the dipoles in the thicker electrolyte provide an additional charge storage facility.)
@@
and at a $5\ nm$ membrane thickness, the voltage across the plates becomes
@@
"The plasma membrane of all cells, including nerve cells, is approximately 6 to 8 nm thick and consists of a mosaic of lipids and proteins." [1], page 71. Using such a value for the membrane’s thickness may result in a value that is up to 60% higher. Furthermore, the usual concentration is also up to 10% higher. Hodgkin, in 1964, measured molarity values in squid axons for ions $K+$, $Na^{+}$, and $Cl^{-}$, (400,50,40-150) inside and (20,440,560) outside, and they provided potential values $55-75~{}mV$ [25]. Given that we used a plausible but ad-hoc "charged layer thickness" and gap distance, we cannot expect a better agreement. However, our qualitative conclusions remain valid and can be adjusted quantitatively to results of dedicated measurements.
Figure 4: The neuronal membrane’s electric field in the function of the distance from the membrane’s surface. The thickness of the ’atomic layers’ proximal to the membrane’s surfaces are also shown. The electric field is linear inside the uniformly charged atomic ion layer on the surface and is described by Eq.(14) in the electrolyte. For visibility, layer thickness $0.5\ nm$ is assumed. For illustration, a simple capped ion channel in the membrane’s wall is also displayed.

### v2

@@ 2.5 One segment
For the discussion below, we assume that the segment has a two-dimensional surface boundary, and we discuss the gradients along a line perpendicular to that plane’s surface. We neglect the electrokinetic and surface-related electric effects connected to membranes and nanopores [19] and focus on the balanced state of membranes in ionic solutions. We compose the segments from such layers (sheets), $(x,y)$ parallel plates, and describe the gradients in the direction of $z$. We introduce the concept of a ’thin physical layer’ parallel to the membrane, characterized by a finite width.
@@
where $N_{A}$ is the number of ions in a millimolmillimol and $e_{el}$ is the elementary charge. Notice that all ions contribute to the charge density, so simply $c=\sum_{k}C_{k}$.
@@ 2.6 Two segments
The behavior near the boundary layers is problematic, primarily because the layers themselves are ill-defined. Physically, they are not ideally plain and well-defined surfaces; the layer thicknesses are limited by the size of the ions/molecules and their associated ions; the layers do not necessarily satisfy the ideal requirements of the respective physical laws. Mathematically, a single ’universal’ function that spans all regions surely cannot accurately describe the behavior of the physical model. So, we chose a piecewise description where the pieces seamlessly fit at the boundaries. We assume that between the surfaces of the membrane, classical electricity describes the electric field; the charged layer and the electrolyte jointly produce the electric field; we have a uniformly charged layer from separated charges on the surface and a polarized electrolyte outside them.
Unlike in the classical theory, the electrolyte segment next to the surface layer contains dipole molecules that have more or less balanced charges (the polarization also depends on the external electric field), so they have much less mobility than the dissociated ions: their size and mass are bigger, and their electric force is a fragment of that of the ions. Their virtual charge becomes apparent only in the appropriate environment, yet it generates an electric field similar to that of a real charge. The final reasons for creating virtual charges are real charges and/or external fields, so an electric field due to virtual charges can also exist inside dielectric matter. We assume the membrane is transparent for the electrical interaction (the electrical field affects the ions in the other segment on the other side of the membrane) but not for their masses (mechanically separates the segments). The membrane does not need to generate any counterforce (except for collisions).
@@ 2.6.1 Charge layers in segments
We consider that the segment is composed of electrically conducting discs (the ions are free to move on the surface), and the charged discs’ contribution to the electric field at point $P$ (see Fig.  2) can be calculated as known from the theory of electricity. Due to the symmetry in the $z$ direction, in a homogeneous solution, the resulting electric field in the plane perpendicular to the $z$ direction is zero, as we have contributions of equal magnitude with opposite signs.
@@ 2.6.2 Simple conducting sheet (a physical condenser)
First, we consider a condenser that obeys the laws of classical electricity and has the geometrical size of a biological neuron. We apply the physics terms to that pseudo-biological object and test whether the values we derive are consistent with physiological phenomena. A tiny (according to [30], page 12, about $2*10^{-3}$) portion of the positive ions leaves their negative counterions behind and forms a thin positively charged ion layer on the surface of the membrane (of course, the same happens on the opposite side with negative ions). Experience shows that – also in biology – two very thin layers of ions are formed on the two sides of the membrane of finite width, see the caption of figure (Fig. 11.22 in [45]): "The ions that give rise to the membrane potential lie in a thin (<1nm<1\ nm) surface layer close to the membrane" (As shown below, these ions give rise only to part of the potential, and the dipoles in the electrolyte contribute the rest). A sub-nanometer-thick conducting layer covers the two surfaces on the top of a 5nm5~nm thick insulator. So, we model the cell with two finite-width "conductive sheet" layers. These layers represent an electric condenser (so we can calculate the internal electric field between the plates) and counterbalance each other’s electric field outside the condenser. We may assume that the solid surface represents the needed counterforce to keep the charges at rest.
We cannot expect geometrically flat surfaces, as some biological objects of up to 1nm1~nm in size are present on the surface. So, we must assume a physical ’uniformly charged sheet’ with an average thickness $\Delta z=0.5~nm$ on the surface. For the sake of simplicity, we assume a step-like function for the electric field. It is zero in the segment outside the "conducting sheets" (the ’bulk’ portions) and inside the sheets. Between the sheets, it jumps to the value of the electric field of a charged sheet with surface density $\sigma_{A}$ (see Eq.(9)). (In this picture, we see a jump in the value of the electric field on the two sides of the physical layer. We assume that the physical layer somehow emulates the well-defined boundary on the side opposite to that proximal to the membrane. Here, we still assume that a mechanical counterforce due to the surface’s roughness keeps the charges inside the charged layer.)
We assume that all unbalanced (dissociated) ions are in the layer, and the rest contains no dissociated ions; furthermore, the concentration of the ions is unchanged in the segment. In the case of a 400mM400~mM solution and a 0.5nm0.5~nm ’thin physical layer’ Eq.(10) evaluates to an electric field
@@
and at a 5nm5\ nm membrane thickness, the voltage across the plates becomes
@@
Our simple model seems to be a strong oversimplification. The number of uncompensated ions needed for the cell, using the method in [30], is $0.88*10^{7}$. The calculated values are about a factor of $5$ lower than the experimentally derived values. "An electrical potential difference about 50100mV50-100\ mV … exists across a plasma membrane only about 5nm5\ nm thick, so that the resulting voltage gradient is about 100, [^0] V/cm100,000\ V/cm[45]. "The number of uncompensated ions needed for the cell is $4.7*10^{7}$[30]. We are in the correct order of magnitude, but we arbitrarily assumed a layer thickness, a non-dielectric medium, and a physically unreasonable step-like electric field. The deviation from the experienced values suggests that dipoles in the electrolyte bulk significantly change the achievable electric field and potential across the plates.
Figure 3: The oppositely charged parallel surfaces of the membrane are treated like conducting plates (infinite planes, neglecting fringing), then one can use Gauss’s law to calculate the electric field between the plates. Presuming the plates are at equilibrium with zero electric fields inside the conductors, the result from a charged conducting surface can be used. “from Hyperphysics by Rod Nave, Georgia State University". ©http://hyperphysics.phy-astr.gsu.edu/hbase/
A usual parallel plate condenser (shown in Fig. 3) can be derived with such a model. We have two infinite conducting sheets and an insulator layer between them. The magnitude of the opposite charges on the two plates must be the same as in the classical picture. Given that the amount of surface charge densities is concluded from the ionic concentration in the segment (in a balanced state), the sum of ionic concentrations must be the same in the segments. A constant electric field is present inside the membrane (across the plates of the condenser), and there is zero electric field inside the parallel conducting plates and outside the condenser. In this ideal picture, the charges are aligned on the border of two (infinitely thin) conducting layers and cannot move. The attractive force between them and the opposite charges on the other plate keeps them fixed in the direction of $z$, and the repulsive force between the charges with the same sign, the surface density in the $(x,y)$ plane remains uniform: the infinitely thin plates are equipotential. This picture illustrates the equilibrium state of charges, consisting of infinitely small non-dissociating charge carriers and perfectly smooth surfaces.
@@ 2.6.3 Condenser in dielectric material
When explaining the effect of dielectricity, we must explicitly consider the duality of ions that they obey laws of electricity and thermodynamics simultaneously; furthermore, the complexity of the electric structure of the solution and that the charge carriers have finite size (see "electron size" vs. "dipole size"). According to the theory of electricity, the free charge carriers (the dissociated ions) are located on the surfaces where they form powerfully charged thin layers (condenser plates) in the segments separated by a membrane on the two surfaces of the membrane; two proximal charge layers on the surfaces of the membrane. We assume that those ions behave as point charges, i.e., due to the attraction from the opposite charges on the opposite plate, they do not produce an electric field toward the side of the bulk of the electrolyte layers. In the classical picture, those layers represent step-like gradients in the electric field along the $z$ axis, and the constraint that the ions must not enter the membrane provides the needed counterforce.
@@
From a thermodynamic point of view, a driving force acts as long as a concentration gradient between neighboring layers exists. From an electrical point of view, the layers are simple parallel-plate condensers that produce no electric field outside their closed volume and are connected serially; plus, the top layer has free ions. The electric field is proportional to the bulk concentration in the layer. Changes in the local electrical gradient may also alter the degree of dissociation and polarization, producing a graded local electric field. These two driving forces have opposite directions and in a balanced state, the same magnitude. The counterforce acts in a way that constrains the molecules to stay in their layers; it adapts (also compensates for the different interaction speeds) while the gradients are changing. It presses the proximal layer to the membrane and the neighboring layers to each other.
@@
Due to the membrane’s finite width and surface roughness, the conducting layer has a finite width. The electric field is homogeneous within the layers. The rest of that electric field toward the bulk layer directs the dipoles proximal to the neighboring dipole layer, and their polarization increases; in other words, they "produce" an electric field. The final effect is that a dipole layer is attracted to the charged layer, and it forms another layer that shows a charged layer toward the bulk. The process repeats and results in a decaying electric field based on the distance from the membrane’s surface.
@@
In our model, we build the volume of the electrolyte from thin layers of directed dipole molecules (the size of molecules limits the thinness $\Delta z$) in the volume, separated into two segments with electrolytes by a membrane with a finite thickness $d$ (that is, a point’s distance in the electrolyte from the two disks will be $z$ and $z+d$, respectively). The contributing potential of an infinitesimal volume on the axis at point $z$ due to the charged sheet is
@@
$U_{left}(c,\Delta z,d)=E^{Gap}_{Classic}(c,\Delta z)*d\int_{-\infty}^{0}\frac{1}{z}dz=\\ E^{Gap}_{Classic}(c,\Delta z)*d\biggl[\ln|z|\biggr]_{0}^{\infty}$ (12)
@@
2.6.4 Leakage current
Although through $\Delta z$, the ion’s water-related behavior can influence the surface charge density, one can conclude that when only one type of ion is solved in the liquid, approximately the same concentration difference between the segments produces approximately the same "thermodynamic electric field" that can counterbalance the electric field of the biological condenser.
HH did a complete measurement, and derived precise measured data. However, their measurement was not accurate because their model was wrong. In a balanced state, no current flows, so no voltage could be generated; see Fig. 3 in [15]: the leakage current is insignificant. More precisely: low intensity slow current flows between nearby ion channels inside the membrane, without involving the AIS. For this reason, they assumed that a permanent leakage current $I_{L}$ was flowing on the membrane, and on its resistance, it generated a 42.5mV42.5\ mV voltage [3]. In HH’s picture, the question what ion constitutes the current, remains open. It was claimed that the only permeability pathways open at rest are $g_{leak}$ and $g_{K}$. It follows that at the resting membrane potential, the leakage current equals the potassium current. That is, at rest, two currents flow (without a driving force in a balanced state), one consisting of $K^{+}$ ions and another of one $leakage^{+/-}$ ions (??) (depending on the direction, it could be positive or negative), one generates voltage while the other not, and neither of them changes the concentration neither on the departure nor on the arrival segment, neither the concentration of the other ions; defying also Nernst’s law. BTW: [47] discusses the operation of the membrane, in some modified HH approach: it replaces "leakage" current with "pump" currents. In the wrong model, such a modification does not matter.
The idea of leakage current was wrong: the voltage HH measured [3] is correct, but – as our model explains – it was generated by charge separation rather than a current through a distributed resistor. It is permanently present while ionic segments are on the two sides of the membrane, even when no current flows. Measuring the energy consumption of neuronal operation [14] confirms that no resting current, in the sense as HH used the concept, exists. The existence of leakage current would be against, among others, the overall metabolic efficiency of neuronal operation.
@@ 2.7 Bridging the micro- and macro views
We assumed that the function can be interpreted for regions $(-\infty,0)$ and $(1,+\infty)$. However, we have arrived at the boundary between quantized and continuous electricity. We use a trick similar to the one Boltzmann used in his famous equation, except that we calculate the number of charge carriers from geometrical instead of statistical assumptions. The minimal thickness of layers is limited, at least by the size of the ion. Furthermore, the membrane’s surface is not flat; there are structures (mainly lipids) with sizes of up to 1nm1~nm, so it is likely realistic to consider a layer thickness of 0.5nm0.5~nm as we did above (on our mathematical scale $0.1=\frac{0.5~nm}{5~nm}$), for which a multiplier of $1.7$ between the classic and the electrolytic fields exists. A similar calculation using a layer thickness of 0.1nm0.1~nm (on our mathematical scale, 0.02) results in a multiplier of $3.2$; so we use a multiplier of value 3 (that corresponds to a layer thickness of 0.125nm0.125~nm (the radius of a $Na^{+}$ ion is 0.116nm0.116~nm)). Correspondingly, we assume an electric field
@@
due to the dielectric properties of the segment. This contribution is due to the dielectricity in the segment (for the sake of simplicity, we have neglected the relative permittivity). It is to be added to the value of the classic contribution, the field generated by the conducting plates, so the final multiplier is $4$. (The experience is known in technical electricity: the electrolytic capacitors achieve several times higher charge storage capacity by using (pseudocapacitance). By using "roughened anode foil", the thickness of the electrolyte is increased, and the roughening provides the necessary mechanical support. A thicker electrolyte layer wraps the condenser plates, and the dipoles in the thicker electrolyte provide an additional charge storage facility.)
@@
and at a 5nm5\ nm membrane thickness, the voltage across the plates becomes
@@
"The plasma membrane of all cells, including nerve cells, is approximately 6 to 8 nm thick and consists of a mosaic of lipids and proteins." [1], page 71. Using such a value for the membrane’s thickness may result in a value that is up to 60% higher. Furthermore, the usual concentration is also up to 10% higher. Hodgkin, in 1964, measured molarity values in squid axons for ions $K^{+}$, $Na^{+}$, and $Cl^{-}$, (400,50,40-150) inside and (20,440,560) outside, and they provided potential values 5575mV55-75~mV [30]. Given that we used a plausible but ad hoc "charged layer thickness" and gap distance, we cannot expect a better agreement. However, our qualitative conclusions remain valid and can be adjusted quantitatively to the results of dedicated measurements.
Figure 4: The neuronal membrane’s electric field in the function of the distance from the membrane’s surface. The thickness of the ’atomic layers’ proximal to the membrane’s surfaces are also shown. The electric field is linear inside the uniformly charged atomic ion layer on the surface and is described by Eq.(14) in the electrolyte. For visibility, a layer thickness 0.5nm0.5\ nm is assumed. For illustration, a simple capped ion channel in the membrane’s wall is also displayed.
@@
2.8 The dynamic layers
This layer’s importance and operation are in line with the Earth’s atmosphere. Its features drastically deviate from the features of the bulks on its two sides. It is separated by a sharp contour on one side and an ill-defined border on the other; its volume is far from being homogeneous. Gravity keeps it in place, and it is at rest. However, sometimes, for some periods, also other (thermodynamical and electrical) forces are evoked inside it and lead to transient changes, moving huge masses with high speeds inside it. Their thickness and mass are negligible compared to those of the bulks on their two sides, and we can describe the bulks without considering their density, mass, size, etc. Still, this thin layer is responsible for the weather. It can temporarily absorb products of slow processes (water evaporation) and deliver masses of high density (much above its density, such as water, sand, etc.) to continental distances, creating the illusion that it stores that matter. Minor changes (natural ones, such as a slight difference in air temperature, and artificial ones, such as injecting condensation nuclei in clouds) can result in enormous changes. We can even imagine volcanic eruptions as semipermeable gates for material with apparently random operation and distribution of the injected material.
To describe those complex and continuous phenomena at least approximately, we must separate them into stages. We can describe the stages approximately using omissions, approximations, and abstractions, usually considering only one dominant phenomenon. The described phenomena are interrelated in a very complex way and depend on different parameters. To some point, we can describe that thin layer using a static picture and provide an empirical description of its processes, even though we can give some limited validity mathematical descriptions for those stages. However, we understand that for describing the time course of the transition (contrasting with step-like stage changes) between those well-defined stages of the atmosphere, we need a dynamic description to discover the laws of motion governing the processes.
Similar is the case with the neuronal membranes and the neuronal operation. Now, we are at the point where their decades-old static description is insufficient. We need to derive the corresponding laws of motion to describe the neuron’s dynamic behavior. We need a meticulous and unusual analysis to derive them. In a neuron, in the abstraction science uses, we put together only an ionic solution, a semipermeable membrane, and currents that enter and leave it. All these belong to non-living matter. As experienced, at some combination of their parameters and gradients, qualitatively different phenomena happen, which, in the abstraction biology uses, are called signs of life; our system starts to belong to living matter. Given that the approximations, the derived abstractions, and the mathematical formalisms describing them are different for the two cases, it looks like we have two different, only loosely bound worlds. We realize we have arrived at the boundary of non-living and living matters, and we must go back to the first principles of science to clarify where their boundary is. However, by using our approach, we may defy that "the emergence of life cannot be predicted by the laws of physics" [48]. Our artificial duck looks, quacks, and swims like a duck.